OpenStudy (anonymous):

4-w/w-6 - 6w-7/6-w (subtract)

OpenStudy (mertsj):

\[\frac{4-w}{w-6}+\frac{6w-7}{w-6}\]

OpenStudy (mertsj):

\[\frac{5w-3}{w-6}\]

OpenStudy (anonymous):

sorry this sight keeps booting me off

OpenStudy (anonymous):

i had a different answer. let me re-work it and see where i made a mistake.

OpenStudy (anonymous):

ok i see what i did. thanks.